Depreciation by Three Methods; Partial Years

Perdue Company purchased equipment on April 1 for $33,210. The equipment was expected to have a useful life of three years, or 3,780 operating hours, and a residual value of $1,080. The equipment was used for 700 hours during Year 1, 1,300 hours in Year 2, 1,100 hours in Year 3, and 680 hours in Year 4.

Required:

Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) units-of-output method, and (c) the double-declining-balance method.

Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar.

a. Straight-line method
Answer #1

a)Straight Line Depreciation Method:

Annual depreciation = (Initial cost – Salvage value)/Useful life

                                    = ($ 33,210 - $ 1,080)/3 = $ 32,130/3 = $ 10,710

Year 1st depreciation = (Number of Months used/12) x Annual depreciation

                                    = 9/12 x $ 10,710

                                    = 0.75 x $ 10,710 = $ 8,032.50

Year 2nd depreciation = $ 10,710

Year 3rd depreciation = $ 10,710

Year 4th depreciation = [(12 - Number of Months used in year 1)/12] x Annual depreciation

                                      = [(12 - 9)/12] x $ 10,710

                                      = 3/12 x $ 10,710 = 0.25 x $ 10,710 = $ 2,677.50

b)Units-of-output Depreciation Method:

Annual depreciation = Depreciable value/ (Number of hours used/estimated operating hours)

                             = (Initial cost – Salvage value)/ (Number of hours used/estimated operating hours)

Year 1st depreciation = ($ 33,210 - $ 1,080) x (700/3,780) = $ 32,130 x 0.185185185185185 = $ 5,950

Year 2nddepreciation = ($ 33,210 - $ 1,080) x (1,300/3,780) = $ 32,130 x 0.343915343915344 =

$ 11,050

Year 3rddepreciation = ($ 33,210 - $ 1,080) x (1,100/3,780) = $ 32,130 x 0.291005291005291= $ 9,350

Year 4thdepreciation = ($ 33,210 - $ 1,080) x (680/3,780) = $ 32,130 x 0.179894179894180 = $ 5,780

c)Double-Decline-Balance Method:

Straight line depreciation percentage = 100%/Useful life = 100%/3 = 33.3333%

Depreciation rate = 2 x Straight line depreciation percentage = 2 x 33.3333% = 66.6666%

Year 1st depreciation rate = No. of months/12 x Depreciation rate

                                    = 9/12 x 66.6666% = 0.75 x 66.6666% = 0.4999995 or 50 %

Year 1st depreciation = Depreciation rate x Initial cost

                                    = 50 % x $ 33,210 = $ 16,605

Year 2nd depreciation = Depreciation rate x Book value at the beginning of period

                                      = Depreciation rate x (Initial cost – Accumulated depreciation)

                                      = 66.6666 % x ($ 33,210 - $ 16,605)

                                      = 66.6666 % x $ 16,605 = $ 11,069.9889 or $ 11,070

Year 3rd depreciation = Depreciation rate x Book value at the beginning of period

                                      = Depreciation rate x (Initial cost – Accumulated depreciation)

                                      = 66.6666 % x ($ 33,210 - $ 16,605 - $ 11,070)

                                      = 66.6666 % x $ 5,535 = $ 3690.0037 or $ 3,690

Year 4th depreciation rate = [(12 - No. of months in year 1st)/12] x Depreciation rate

                                              = (12-9)/12 x 66.6666 %

                                              = 3/12 x 66.6666 % = 0.25 x 66.6666 % = 16.6667 %

Year 4th depreciation = Depreciation rate x Book value at the beginning of period

                                      = 16.6667 % x ($ 33,210 - $ 16,605 - $ 11,070 - $ 3,690)

                                      = 16.6667 % x $ 1,845 = $ 307.5006 or $ 308